EDITORIAL

Terry Goodman

It seems only a short time ago that the first issue of MJMS was published. With this current issue we begin our fifth year of publication. Over the previous four years, MJMS has evolved in a variety of ways — format, emphases, etc.

We want to make you aware of a change that we are making in our editing and reviewing process. We will be using five Associate Editors, each of whom will be responsible for editing articles from specific areas of mathematics. In the future, manuscripts should be sent to these Associate Editors to begin the reviewing process. We believe that this will help to streamline this process and minimize the time from submission to disposition of a manuscript. We also like the idea of more people being directly involved with MJMS.

We welcome any comments or suggestions you have concerning these changes. Below is a list of the Associate Editors and the areas for which they will be responsible.

Analysis and Differential Equations Computer Science & Applied Mathematics

Dr. Leon Hall  Mr. William Grimes
Department of Mathematics  Department of Mathematics
& Statistics  & Computer Science
University of Missouri–Rolla  Central Missouri State University
Rolla, MO 65401-0249  Warrensburg, MO 64093

Mathematics Education Number Theory, Algebra, & Misc.

Dr. David Ewing  Dr. Robert Kennedy
Department of Mathematics  Department of Mathematics
& Computer Science  & Computer Science
Central Missouri State University  Central Missouri State University
Warrensburg, MO 64093  Warrensburg, MO 64093

Topology and Geometry

Dr. Shing So Department of Mathematics

& Computer Science Central Missouri State University Warrensburg, MO 64093